If the total cost of producing three units of output is $2,621 and the total cost of producing four units of output is $3,109 then the marginal cost of the fourth unit is? Please enter your response as an integer.
Answer #1

Answer

Marginal cost is the additional cost incurred in order to produce one additional unit of output.

Mathematically, MC(Q) = TC(Q) - TC(Q-1)

where MC(Q) = Marginal cost of producing Qth unit and TC(Q) = Total cost of producing Q units

Here we have to calculate MC(4).

using above formula :

MC(4) = TC(4) - TC(3) = 3109 - 2621 = 488

Hence, the marginal cost of the fourth unit is 488.